OpenStudy (anonymous):

there are 1400 workers in the factory. 5/7 of them are men and 3/5 of the men are greek. How many greek men work in the factory?

OpenStudy (lgbasallote):

5/7 of 1400 is 5/7 x 1400 = 5 x 200 = 1000 3/5 of that is greek so 3/5 x 1000 what do you get?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 1400 \cdot \large {5 \over 7} \cdot {3 \over 5} }\) Simplicity ^_^

Parth (parthkohli):

'of' ===> MULTIPLY

Parth (parthkohli):

3/5 OF 5/7 OF 1400

OpenStudy (anonymous):

Solution is : 3/5 of 5/7 of 1400 = Whatever you get...